OpenStudy (anonymous):

A shop sells compact disks for $18 per piece. If each disk costs $10 to the shop owner how much profit percent does the shop owner earn?

OpenStudy (anonymous):

\[(18-10)/18 = 4/9 = 0.444...\]

OpenStudy (anonymous):

er, 44.4%

OpenStudy (anonymous):

thanks got that also. yay!